OpenStudy (ghostedly):

A wheelchair ramp starts 8 meters from the base of a staircase. The staircase is 1 meter high. What is the length of the wheelchair ramp? Round your answer to the nearest tenth.

OpenStudy (ghostedly):

@Dani_Rose

OpenStudy (impracticaljoker):

Use slope intercept form

OpenStudy (dani_rose):

i don't think you use slope intercept form

OpenStudy (impracticaljoker):

Im so sorry. Wrong question. On this one you use pythagarem theorem which is a^2 + b ^2 = c^2 and I know I spelled pythagarem wrong

OpenStudy (dani_rose):

Formula: a^2 + b^2 = c^2. 8^2 + 1^2 = c^2 64 + 1 = c^2 c^2 = 65 Square root of 65 is 8.06 which rounds to 8.1